### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) has emerged as a pivotal innovation in natural language processing, enhancing generative models by incorporating external information retrieval. Evaluating RAG systems, however, poses distinct challenges due to their hybrid structure and reliance on dynamic knowledge sources. We consequently enhanced an extensive survey and proposed an analysis framework for benchmarks of RAG systems, RGAR (Retrieval, Generation, Additional Requirement), designed to systematically analyze RAG benchmarks by focusing on measurable outputs and established truths. Specifically, we scrutinize and contrast multiple quantifiable metrics of the Retrieval and Generation component, such as relevance, accuracy, and faithfulness, of the internal links within the current RAG evaluation methods, covering the possible output and ground truth pairs. We also analyze the integration of additional requirements of different works, discuss the limitations of current benchmarks, and propose potential directions for further research to address these shortcomings and advance the field of RAG evaluation. In conclusion, this paper collates the challenges associated with RAG evaluation. It presents a thorough analysis and examination of existing methodologies for RAG benchmark design based on the proposed RGAR framework.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) [29] represents a significant advancement in natural language processing by enhancing the performance of generative models through integrating information retrieval techniques. It addresses a critical challenge faced by standalone generative models: the tendency to produce responses that, while plausible, may not be grounded in facts. By retrieving relevant information from external sources, RAG significantly reduces the incidence of hallucinations [20] or factually incorrect generative outputs, thereby improving the content’s reliability and richness. [56] This fusion of retrieval and generation capabilities enables the creation of responses that are not only contextually appropriate but also informed by the most current and accurate information available, making RAG a development in the pursuit of more intelligent and versatile language models [56, 52].
Figure 1: The structure of the RAG system with retrieval and generation components and corresponding four phrases: indexing, search, prompting and inferencing. The pairs of EOs and GTs are highlighted in red and green, with brown dashed arrows.
@@
Numerous studies of RAG systems have emerged from various perspectives since the advent of pre-trained language models [14]. The RAG system comprises two primary components: Retrieval and Generation. The retrieval component aims to extract relevant information from a vast array of external knowledge sources. It involves two main phases, indexing and searching. Indexing organizes documents to facilitate efficient retrieval, using either inverted indexes for sparse retrieval or dense vector encoding for dense retrieval [14, 11, 24]. The searching utilizes these indexes to fetch relevant documents on the user’s query, often incorporating the optional rerankers [4, 34, 5, 43] to refine the ranking of the retrieved documents. The generation component utilizes the retrieved content to formulate coherent and contextually relevant responses with the prompting and inferencing phases. The input for language models is formulated through prompting, integrating the query from the retrieval phase. Methods like Chain of Thought (CoT) [48] or Rephrase and Respond (RaR) [7] guide better generation results. In the inferencing step, Large Language Models (LLMs) interpret the prompted input to generate accurate and in-depth responses that align with the query’s intent and integrate the extracted information [30, 8]. The RAG structure is explained in detail in Appendix 0.A. Figure 1 illustrates the structure of the RAG systems as mentioned.
@@
The significance of evaluating RAG is escalating in tandem with the advancement of RAG-specific methodologies. On the one hand, RAG is a complex system intricately tied to specific requirements and language models, resulting in a variety of evaluation methods, indicators, and tools, particularly given the black-box LLM generation. Evaluating RAG systems thus involves considering quite a few specific components and the complexity of overall system assessment. On the other hand, the complexity of RAG systems is further compounded by the dynamic external database and the various downstream tasks, such as content creation or open domain question answering [14, 53]. These challenges necessitate the development of comprehensive evaluation metrics that can effectively capture the interplay between retrieval accuracy and generative quality [2, 6]. To clarify the elements further, we conducted this survey on RAG evaluation to address the current gaps in the area, which differs from the prior RAG surveys [57, 14, 21] that predominantly collected specific RAG methods or data. We have compiled 12 distinct evaluation frameworks, encompassing a range of aspects of the RAG system. We conduct a comparative analysis and synthesize the specific evaluation methods of various components, focusing on aspects such as accuracy, faithfulness, and relevance. We also discuss the constraints of the existing methodology and the prospects for future RAG evaluations. We hope to provide the readers with a comprehensive understanding of the RAG evaluation.
@@
    Challenge of Evaluation: This is the first work that summarize and classify the challenges in evaluating RAG systems through the structure of RAG systems, including three parts retrieval, generation, and the whole system.
@@
    Analysis Framework: Based on the challenges, we propose an analysis framework (RGAG) for RAG benchmarks, which is designed to navigate the unique complexities inherent to RAG systems, offering a fundamental methodology for assessing their efficacy across many facets.
@@
    RAG Benchmark Analysis: With the help of the RGAG framework, we provide a comprehensive analysis of existing RAG benchmarks, highlighting their strengths and limitations and proposing recommendations for future developments in RAG system evaluation.
@@ 2 Challenges in Evaluating RAG Systems
The retrieval component of RAG systems is critical for fetching relevant information from external knowledge sources that inform the generation process. One primary challenge in evaluating this component is the dynamic and vast nature of potential knowledge bases, ranging from structured databases to the entire web. This vastness requires evaluation metrics that can effectively measure the precision, recall, and relevance of retrieved documents in the context of a given query [43, 27]. Moreover, the temporal aspect of information, where the relevance and accuracy of data can change over time, adds another layer of complexity to the evaluation process [5]. Additionally, the diversity of information sources and the potential for retrieving misleading or low-quality information poses significant challenges in assessing the retrieval component’s effectiveness in filtering and selecting the most pertinent information [34]. On the other hand, the evaluation method of the retrieval component is also a challenge, as the traditional evaluation metrics enforce the retrieval system focus on higher TopK recall, instead of the useful information carried through one query.
@@
The generation component, currently powered by LLMs, produces coherent and contextually appropriate responses based on the retrieved content. A significant challenge here lies in evaluating the faithfulness and accuracy of the generated content to the input data. This involves not only assessing the factual correctness of responses but also their relevance to the original query and the coherence of the generated text [58, 41]. The subjective nature of certain tasks, such as creative content generation or open-ended question answering, further complicates the evaluation, as it introduces variability in what constitutes a ‘correct’ or ‘high-quality’ response [40].
@@
Evaluating the whole RAG system introduces additional complexities. The interplay between the retrieval and generation components means that the entire system’s performance cannot be fully understood by evaluating each component in isolation [41, 13]. The system needs to be assessed on its ability to leverage retrieved information effectively to improve response quality, which involves measuring the added value of the retrieval component to the generative process. Furthermore, practical considerations such as response latency, robustness against misinformation, and the ability to handle ambiguous or complex queries are also crucial for evaluating the system’s overall effectiveness and usability [34, 5].
3 RGAR: Analysis Framework for Evaluation
Evaluating the target shift from traditional absolute numeric metrics to multi-source and multi-target generation evaluation, along with the intricate interplay between retrieval and generation components, poses significant challenges. Searches in a dynamic database may lead to misleading results or contradicting the facts. Diverse and comprehensive datasets that accurately reflect real-world scenarios are crucial. Challenges also arise in the realm of metrics, encompassing generative evaluation criteria for distinct downstream tasks, human preferences, and practical considerations within the RAG system. Most prior benchmarks predominantly tackle one or several aspects of the RAG assessment but lack a comprehensive, holistic analysis.
To provide a better understanding of RAG benchmarks, we propose an analysis framework named RGAR (Retrieval, Generation, and Additional Requirement). It takes into account the Target, Dataset, and Metric respectively. The Target module is intended to determine the evaluation direction. The Dataset module facilitates the comparison of various data constructions in RAG benchmarks. The final module, Metrics, introduces the metrics that correspond to specific targets and datasets used during evaluation. Overall, it is designed to provide a systematic methodology for assessing the effectiveness of RAG systems across various aspects by covering all possible pairs between the “Evaluable Outputs” (EOs) and “Ground Truths” (GTs). In the following section, we will explain thoroughly the framework and utilize it for introducing and comparing the RAG benchmarks.
Figure 2: The Target modular of RGAR framework.
@@ 3.1 Evaluation Target (What to Evaluate?)
The combination of EOs and GTs in the RAG system can generate all possible targets, which is the fundamental concept of the RGAR framework (as shown in Figure 1). Once identified, these targets can be defined based on a specific pair of EOs or EO with GT, as illustrated in Figure 2, and used to analyze all aspects of current RAG benchmarks.
Retrieval
For the evaluation of the retrieval component, the EOs are the relevant documents depending on the query. Then we can construct two pairwise relationships for the retrieval component, which are Relevant Documents $\leftrightarrow$ Query, Relevant Documents $\leftrightarrow$ Documents Candidates.
@@
Tool TruEra RAG Triad [45] 2023.10
Context Relevance
Answer Relevance
Groundedness
LLM as Judge LLM as Judge
Tool LangChain Bench. [27] 2023.11
Accuracy
Faithfulness
Execution Time
Embed. CosDistance
Accuracy LLM as Judge
Tool Databricks Eval [28] 2023.12
Correctness
Readability
Comprehensiveness
- LLM as Judge
Benchmark RAGAs [13] 2023.09
Context Relevance
Answer Relevance
Faithfulness
LLM as Judge
LLM Gen + CosSim
LLM as Judge
Benchmark RECALL [33] 2023.11
Response Quality
Robustness
- BLEU, ROUGE-L
Benchmark ARES [41] 2023.11
Context Relevance
Answer Faithfulness
Answer Relevance
LLM + Classifier
LLM + Classifier
LLM + Classifier
Benchmark RGB [5] 2023.12
Information Integration
Noise Robustness
Negative Rejection
Counterfactual Robustness
- Accuracy
Benchmark MultiHop-RAG [43] 2024.01
Retrieval Quality
Response Correctness
MAP, MRR, Hit@K LLM as Judge
Benchmark CRUD-RAG [34] 2024.02
CREATE, READ
UPDATE, DELETE
ROUGE, BLEU
RAGQuestEval
Benchmark MedRAG [49] 2024.02
Accuracy
- Accuracy
Benchmark FeB4RAG [47] 2024.02
Consistency
Correctness
Clarity
Coverage
Human Evaluation
Human Evaluation
Benchmark CDQA [50] 2024.03
Accuracy
- F1 of Tokens
@@
Research Diversity Reranker [4] 2023.08 Diversity - -
Generation
@@
The targets of Retrieval and Generation components are introduced. Table 1 lists the relative work about improving and evaluating RAG and RAG benchmarks cut off in March 2024. Table 1 portrays this information, where each evaluation criterion is represented by a different colour. For example, FeB4RAG [47], the fourth from the last, has posited four standards based on [15] that comprise Consistency, Correctness, Clarity, and Coverage. Correctness is equivalent to accuracy in retrieval, and Consistency is tantamount to faithfulness in the generation component. While accuracy in retrieval gauges the correctness of the retrieved information, we posit that Coverage pertains to the coverage rate and is more associated with diversity. Therefore, we consider Coverage to be linked with diversity and an additional requirement in our proposed evaluation framework, which will be introduced subsequently. The remaining standard, Clarity, is also classified as an additional requirement in our proposed framework. The other tools and benchmarks are processed similarly.
Tools and benchmarks offer varying degrees of flexibility in evaluating datasets for RAG systems. Tools, which specify only evaluation targets, provide a versatile framework capable of constructing complete RAG applications and evaluation pipelines, as seen in works like [45, 27, 28]. Benchmarks, on the other hand, focus on different aspects of RAG evaluation with specific emphasis on either retrieval outputs or generation targets. For instance, RAGAs and ARES assess the relevance of retrieval documents, while RGB and MultiHop-RAG [5, 43] prioritize accuracy, necessitating comparison with GTs. All benchmarks consider generation targets due to their critical role in RAG systems, though their focus areas vary.
Additional Requirement
In addition to evaluating the two primary components outlined, a portion of the works also addressed some additional requirements of RAG (Black and Ltatic targets in the Table 2). The requirements are as follows:
@@
For the exception, CRUD-RAG [34] introduces a comprehensive benchmark addressing the broader spectrum of RAG applications beyond question-answering, categorized into Create, Read, Update, and Delete scenarios. This benchmark evaluates RAG systems across diverse tasks, including text continuation, question answering, hallucination modification, and multi-document summarization, each corresponding to CRUD actions. It emphasizes the evaluation of all RAG components, including retrieval models and external knowledge base construction, offering insights for optimizing RAG technology across different scenarios.
3.2 Evaluation Dataset (How to evaluate?)
@@
CDQA [50] Generations (Source: News), Labeller
In Table 2, distinct benchmarks employ varying strategies for dataset construction, ranging from leveraging existing resources to generating entirely new data tailored for specific evaluation aspects. Several benchmarks draw upon the part of KILT (Knowledge Intensive Language Tasks) benchmark [37] (Natural Questions [25], HotpotQA [51], and FEVER [44])and other established datasets such as SuperGLUE [46] (MultiRC [9], and ReCoRD [54]) [41]. However, the drawback of using such datasets can’t solve the challenges in dynamic real-world scenarios. A similar situation can be observed in WikiEval constructed by RAGAs [13], which was generated from Wikipedia pages post 2022.
The advent of powerful LLMs has revolutionized the process of dataset construction. With the ability to design queries and ground truths for specific evaluation targets using these frameworks, authors can now create datasets in the desired format with ease. Benchmarks like RGB, MultiHop-RAG, CRUD-RAG, and CDQA [5, 43, 34, 50] have taken this approach further by building their own datasets using online news articles to test RAG systems’ ability to handle real-world information beyond the training data of LM frameworks.

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) has recently gained traction in natural language processing. Numerous studies and real-world applications are leveraging its ability to enhance generative models through external information retrieval. Evaluating these RAG systems, however, poses unique challenges due to their hybrid structure and reliance on dynamic knowledge sources. To better understand these challenges, we conduct A Unified Evaluation Process of RAG (Auepora) and aim to provide a comprehensive overview of the evaluation and benchmarks of RAG systems. Specifically, we examine and compare several quantifiable metrics of the Retrieval and Generation components, such as relevance, accuracy, and faithfulness, within the current RAG benchmarks, encompassing the possible output and ground truth pairs. We then analyze the various datasets and metrics, discuss the limitations of current benchmarks, and suggest potential directions to advance the field of RAG benchmarks.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) [34] efficiently enhances the performance of generative language models through integrating information retrieval techniques. It addresses a critical challenge faced by standalone generative language models: the tendency to produce responses that, while plausible, may not be grounded in facts. By retrieving relevant information from external sources, RAG significantly reduces the incidence of hallucinations [23] or factually incorrect outputs, thereby improving the content’s reliability and richness. [73] This fusion of retrieval and generation capabilities enables the creation of responses that are not only contextually appropriate but also informed by the most current and accurate information available, making RAG a development in the pursuit of more intelligent and versatile language models [73, 64].
@@
Numerous studies of RAG systems have emerged from various perspectives since the advent of Large Language Models (LLMs) [55, 45, 59, 42, 41, 69, 16]. The RAG system comprises two primary components: Retrieval and Generation. The retrieval component aims to extract relevant information from various external knowledge sources. It involves two main phases, indexing and searching. Indexing organizes documents to facilitate efficient retrieval, using either inverted indexes for sparse retrieval or dense vector encoding for dense retrieval [16, 12, 28]. The searching component utilizes these indexes to fetch relevant documents on the user’s query, often incorporating the optional rerankers [4, 39, 6, 52] to refine the ranking of the retrieved documents. The generation component utilizes the retrieved content and question query to formulate coherent and contextually relevant responses with the prompting and inferencing phases. As the “Emerging” ability [59] of LLMs and the breakthrough in aligning human commands [42], LLMs are the best performance choices model for the generation stage. Prompting methods like Chain of Thought (CoT) [60], Tree of Thgouht [65], Rephrase and Respond (RaR) [8] guide better generation results. In the inferencing step, LLMs interpret the prompted input to generate accurate and in-depth responses that align with the query’s intent and integrate the extracted information [35, 9] without further finetuning, such as fully finetuning [16, 1, 67, 68] or LoRA [21]. Appendix 0.A details the complete RAG structure. Figure 1 illustrates the structure of the RAG systems as mentioned.
Figure 1: The structure of the RAG system with retrieval and generation components and corresponding four phrases: indexing, search, prompting and inferencing. The pairs of “Evaluable Outputs” (EOs) and “Ground Truths” (GTs) are highlighted in read frame and green frame, with brown dashed arrows.
@@
The importance of evaluating RAG is increasing in parallel with the advancement of RAG-specific methodologies. On the one hand, RAG is a complex system intricately tied to specific requirements and language models, resulting in various evaluation methods, indicators, and tools, particularly given the black-box LLM generation. Evaluating RAG systems involves specific components and the complexity of the overall system assessment. On the other hand, the complexity of RAG systems is further compounded by the external dynamic database and the various downstream tasks, such as content creation or open domain question answering [16, 70]. These challenges necessitate the development of comprehensive evaluation metrics that can effectively capture the interplay between retrieval accuracy and generative quality [2, 7]. To clarify the elements further, we try to address the current gaps in the area, which differs from the prior RAG surveys [74, 16, 24] that predominantly collected specific RAG methods or data. We have compiled 12 distinct evaluation frameworks, encompassing a range of aspects of the RAG system. Following the procedure of making benchmarks, we analyze through targets, datasets and metrics mentioned in these benchmarks and summarize them into A Unified Evaluation Process of RAG (Auepora) as three corresponding phases.
@@
    Challenge of Evaluation: This is the first work that summarizes and classifies the challenges in evaluating RAG systems through the structure of RAG systems, including three parts retrieval, generation, and the whole system.
@@
    Analysis Framework: In light of the challenges posed by RAG systems, we introduce an analytical framework, referred to as A Unified Evaluation Process of RAG (Auepora), which aims to elucidate the unique complexities inherent to RAG systems and guide for readers to comprehend the effectiveness of RAG benchmarks across various dimensions
@@
    RAG Benchmark Analysis: With the help of Auepora, we comprehensively analyze existing RAG benchmarks, highlighting their strengths and limitations and proposing recommendations for future developments in RAG system evaluation.
@@ 2 Challenges in Evaluating RAG Systems
The retrieval component is critical for fetching relevant information that informs the generation process. One primary challenge is the dynamic and vast nature of potential knowledge bases, ranging from structured databases to the entire web. This vastness requires evaluation metrics that can effectively measure the precision, recall, and relevance of retrieved documents in the context of a given query [52, 32]. Moreover, the temporal aspect of information, where the relevance and accuracy of data can change over time, adds another layer of complexity to the evaluation process [6]. Additionally, the diversity of information sources and the possibility of retrieving misleading or low-quality information pose significant challenges in assessing the effectiveness of filtering and selecting the most pertinent information [39]. The traditional evaluation indicators for retrieval, such as Recall and Precision, cannot fully capture the nuances of RAG retrieval systems, necessitating the development of more nuanced and task-specific evaluation metrics [49].
@@
The generation component, powered by LLMs, produces coherent and contextually appropriate responses based on the retrieved content. The challenge here lies in evaluating the faithfulness and accuracy of the generated content to the input data. This involves not only assessing the factual correctness of responses but also their relevance to the original query and the coherence of the generated text [75, 49]. The subjective nature of certain tasks, such as creative content generation or open-ended question answering, further complicates the evaluation, as it introduces variability in what constitutes a “correct” or “high-quality” response [48].
@@
Evaluating the whole RAG system introduces additional complexities. The interplay between the retrieval and generation components means that the entire system’s performance cannot be fully understood by evaluating each component in isolation [49, 14]. The system needs to be assessed on its ability to leverage retrieved information effectively to improve response quality, which involves measuring the added value of the retrieval component to the generative process. Furthermore, practical considerations such as response latency and the ability to handle ambiguous or complex queries are also crucial for evaluating the system’s overall effectiveness and usability [39, 6].
Conclusion
Evaluating the target shift from traditional absolute numeric metrics to multi-source and multi-target generation evaluation, along with the intricate interplay between retrieval and generation components, poses significant challenges. [5, 50] Searches in a dynamic database may lead to misleading results or contradict the facts. Diverse and comprehensive datasets that accurately reflect real-world scenarios are crucial. Challenges also arise in the realm of metrics, encompassing generative evaluation criteria for distinct downstream tasks, human preferences, and practical considerations within the RAG system. Most prior benchmarks predominantly tackle one or several aspects of the RAG assessment but lack a comprehensive, holistic analysis.
3 A Unified Evaluation Process of RAG (Auepora)
To facilitate a deeper understanding of RAG benchmarks, we introduce A Unified Evaluation Process of RAG (Auepora), which focuses on three key questions of benchmarks: What to Evaluate? How to Evaluate? How to Measure? which correlated to Target, Dataset, and Metric respectively. We aim to provide a clear and accessible way for readers to comprehend the complexities and nuances of RAG benchmarking.
The Target module is intended to determine the evaluation direction. The Dataset module facilitates the comparison of various data constructions in RAG benchmarks. The final module, Metrics, introduces the metrics that correspond to specific targets and datasets used during evaluation. Overall, it is designed to provide a systematic methodology for assessing the effectiveness of RAG systems across various aspects by covering all possible pairs at the beginning between the “Evaluable Outputs” (EOs) and “Ground Truths” (GTs). In the following section, we will explain thoroughly Aueporaand utilize it for introducing and comparing the RAG benchmarks.
Figure 2: The Target modular of the Auepora.
@@ 3.1 Evaluation Target (What to Evaluate?)
The combination of EOs and GTs in the RAG system can generate all possible targets, which is the fundamental concept of the Auepora (as shown in Figure 1). Once identified, these targets can be defined based on a specific pair of EOs or EO with GT, as illustrated in Figure 2, and used to analyze all aspects of current RAG benchmarks.
3.1.1 Retrieval
The EOs are the relevant documents for evaluating the retrieval component depending on the query. Then we can construct two pairwise relationships for the retrieval component, which are Relevant Documents $\leftrightarrow$ Query, Relevant Documents $\leftrightarrow$ Documents Candidates.
@@ 3.1.1 Retrieval
Tool TruEra RAG Triad [54] 2023.10 Context Relevance Answer Relevance Groundedness LLM as a Judge LLM as a Judge
Tool LangChain Bench. [32] 2023.11 Accuracy Faithfulness Execution Time Embed. CosDistance Accuracy LLM as a Judge
Tool Databricks Eval [33] 2023.12 Correctness Readability Comprehensiveness - LLM as a Judge
Benchmark RAGAs [14] 2023.09 Context Relevance Answer Relevance Faithfulness LLM as a Judge LLM Gen + CosSim LLM as a Judge
Benchmark RECALL [38] 2023.11 Response Quality Robustness - BLEU, ROUGE-L
Benchmark ARES [49] 2023.11 Context Relevance Answer Faithfulness Answer Relevance LLM + Classifier LLM + Classifier LLM + Classifier
Benchmark RGB [6] 2023.12 Information Integration Noise Robustness Negative Rejection Counterfactual Robustness - Accuracy
Benchmark MultiHop-RAG [52] 2024.01 Retrieval Quality Response Correctness MAP, MRR, Hit@K LLM as a Judge
Benchmark CRUD-RAG [39] 2024.02 CREATE, READ UPDATE, DELETE - ROUGE, BLEU RAGQuestEval
Benchmark MedRAG [61] 2024.02 Accuracy - Accuracy
Benchmark FeB4RAG [57] 2024.02 Consistency Correctness Clarity Coverage - Human Evaluation Human Evaluation
Benchmark CDQA [62] 2024.03 Accuracy - F1
Benchmark DomainRAG [58] 2024.06 Correctness Faithfulness Noise Robustness Structural Output - F1, Exact-Match Rouge-L LLM as a Judge
Benchmark ReEval [66] 2024.06 Hallucination - F1, Exacct-Match LLM as a Judge Human Evaluation
@@
Research Diversity Reranker [4] 2023.08 Diversity Cosine Distance -
3.1.2 Generation
@@
The targets of Retrieval and Generation components are introduced. Table 1 lists the relative work on improving and evaluating RAG and its benchmarks cut off in June 2024. Table 1 portrays this information, where each evaluation criterion is represented by a different colour. For example, FeB4RAG [57], the fourth from the last, has posited four standards based on [17] that comprise Consistency, Correctness, Clarity, and Coverage. Correctness is equivalent to accuracy in retrieval, and Consistency is tantamount to faithfulness in the generation component. While accuracy in retrieval gauges the correctness of the retrieved information, we posit that Coverage pertains to the coverage rate and is more associated with diversity. Therefore, we consider Coverage to be linked with diversity and an additional requirement in our proposed evaluation framework, which will be introduced subsequently. The remaining standard, Clarity, is also classified as an additional requirement in our proposed framework. The other tools and benchmarks are processed similarly.
Tools and benchmarks offer varying degrees of flexibility in evaluating datasets for RAG systems. Tools, which specify only evaluation targets, provide a versatile framework capable of constructing complete RAG applications and evaluation pipelines, as seen in works like [54, 32, 33]. Benchmarks, on the other hand, focus on different aspects of RAG evaluation with specific emphasis on either retrieval outputs or generation targets. For instance, RAGAs [14] and ARES [49] assess the relevance of retrieval documents, while RGB and MultiHop-RAG [6, 52] prioritize accuracy, necessitating comparison with GTs. The [66] focuses on the Hallucination, which is a combination of faithfulness and correctness. All benchmarks consider generation targets due to their critical role in RAG systems, though their focus areas vary.
3.1.3 Additional Requirement
In addition to evaluating the two primary components outlined, a portion of the works also addressed some additional requirements of RAG (Black and Italics targets in Table 2). The requirements are as follows:
@@ 3.1.3 Additional Requirement
For the exception, CRUD-RAG [39] introduces a comprehensive benchmark addressing the broader spectrum of RAG applications beyond question-answering, categorized into Create, Read, Update, and Delete scenarios. This benchmark evaluates RAG systems across diverse tasks, including text continuation, question answering, hallucination modification, and multi-document summarization. It offers insights for optimizing RAG technology across different scenarios. DomainRAG [58] identifies six complex abilities for RAG systems: conversational, structural information, faithfulness, denoising, time-sensitive problem solving, and multi-doc understanding. ReEval [66] specifically targets hallucination evaluation by employing a cost-effective LLM-based framework that utilizes prompt chaining to create dynamic test cases.
@@
CDQA [62] Generation (Source: News), Labeller
DomainRAG [58] Generation (Source: College Admission Information)
ReEval [66] RealTimeQA[27], NQ [15, 29])
3.2 Evaluation Dataset (How to evaluate?)
In Table 2, distinct benchmarks employ varying strategies for dataset construction, ranging from leveraging existing resources to generating entirely new data tailored for specific evaluation aspects. Several benchmarks draw upon the part of KILT (Knowledge Intensive Language Tasks) benchmark [44] (Natural Questions [29], HotpotQA [63], and FEVER [53]) and other established datasets such as SuperGLUE [56] (MultiRC [10], and ReCoRD [71]) [49]. However, the drawback of using such datasets can’t solve the challenges in dynamic real-world scenarios. A similar situation can be observed in WikiEval, from Wikipedia pages post 2022, constructed by RAGAs [14].
The advent of powerful LLMs has revolutionized the process of dataset construction. With the ability to design queries and ground truths for specific evaluation targets using these frameworks, authors can now create datasets in the desired format with ease. Benchmarks like RGB, MultiHop-RAG, CRUD-RAG, and CDQA [6, 52, 39, 62] have taken this approach further by building their own datasets using online news articles to test RAG systems’ ability to handle real-world information beyond the training data of LM frameworks. Most recently, DomainRAG [58] combines various types of QA datasets with single-doc, multi-doc, single-round, and multi-round. These datasets are generated from the yearly changed information from the college website for admission and enrollment, which forces the LLMs to use the provided and updated information.